# Towards High-Quality and Efficient Speech Bandwidth Extension with Parallel Amplitude and Phase Prediction

### v1

@@ Abstract
Speech bandwidth extension (BWE) refers to widening the frequency bandwidth range of speech signals, enhancing the speech quality towards brighter and fuller. This paper proposes a generative adversarial network (GAN) based BWE model with parallel prediction of Amplitude and Phase spectra, named AP-BWE, which achieves both high-quality and efficient wideband speech waveform generation. The proposed AP-BWE generator is entirely based on convolutional neural networks (CNNs). It features a dual-stream architecture with mutual interaction, where the amplitude stream and the phase stream communicate with each other and respectively extend the high-frequency components from the input narrowband amplitude and phase spectra. To improve the naturalness of the extended speech signals, we employ a multi-period discriminator at the waveform level and design a pair of multi-resolution amplitude and phase discriminators at the spectral level, respectively. Experimental results demonstrate that our proposed AP-BWE achieves state-of-the-art performance in terms of speech quality for BWE tasks targeting sampling rates of both 16 kHz and 48 kHz. In terms of generation efficiency, due to the all-convolutional architecture and all-frame-level operations, the proposed AP-BWE can generate 48 kHz waveform samples 292.3 times faster than real-time on a single RTX 4090 GPU and 18.1 times faster than real-time on a single CPU. Notably, to our knowledge, AP-BWE is the first to achieve the direct extension of the high-frequency phase spectrum, which is beneficial for improving the effectiveness of existing BWE methods.
@@ I Introduction
In the time domain, speech BWE is a conditionally stringent form of speech super-resolution (SR). Speech SR aims to increase the temporal resolution of low-resolution speech signals by generating high-frequency components, where low-resolution speech signals may contain aliased high-frequency components. In contrast, in BWE, only the low-frequency components are preserved in the narrowband signals. Consequently, the BWE task poses greater challenges than SR. Nevertheless, the majority of SR methods are applicable to the BWE task.
@@ III-A1 Generator
where $\bm{\hat{Y}}_{r}=\bm{\hat{Y}}_{a}\cdot\cos(\bm{\hat{Y}}_{p})\in\mathbb{R}^{T\times F}$ and $\bm{\hat{Y}}_{i}=\bm{\hat{Y}}_{a}\cdot\sin(\bm{\hat{Y}}_{p})\in\mathbb{R}^{T\times F}$ denote the real and imaginary parts of the extended short time complex spectrum, respectively.
@@ III-A2 Discriminator
Directly predicting amplitude and phase and then reconstructing the speech waveform through iSTFT can result in over-smoothed spectral parameters, manifesting as a robotic or muffled quality in the reconstructed waveforms. To this end, we utilize discriminators defined both in the spectral domain and time domain to guide the generator $G$ in generating spectra and waveforms that closely resemble real ones. Firstly, considering that the speech signal is composed of sinusoidal signals with various frequencies, with some frequency bands generated through BWE. Due to the statistical characteristics of speech signals varying in different frequency bands, we employ an MPD [40] to capture periodic patterns, with the aim of matching the natural wideband speech across multiple frequency bands. Moreover, since the statistical characteristics of amplitude and phase also differ across frequency bands, and the sole utilization of MPD cannot cover all frequency bands, we consequently define discriminators on both amplitude and phase spectra. Drawing inspiration from the multi-resolution discriminator [41], we respectively introduce MRAD and MRPD, with the aim to capture full-band amplitude and phase patterns at various resolutions. The details of MPD, MRAD, and MRPD are described as follows.
@@ III-B3 Final Loss
$\mathcal{L}_{G}=\sum_{k=1}^{K}\bigg{[}\mathcal{L}_{adv}(G;D_{k})+\mathcal{L}_{FM}(G;D_{k})\bigg{]}+\mathcal{L}_{S},$ (13)
@@
where $K$ denotes the numbers of sub-discriminators, and $D_{k}$ denotes the $k$-th sub-discriminator in MPD, MRAD, and MRPD.
@@ IV-B Model Details
We used the same configuration for experiments with target sampling rates of 16 kHz and 48 kHz. For training our proposed AP-BWE model, all the audio clips underwent silence trimming with VCTK silence labels  [^1] https://github.com/nii-yamagishilab/vctk-silence-labels. and sliced into 8000-sample-point segments. To extract the amplitude and phase spectra from raw waveforms, we used STFT with the FFT point number of 1024, Hanning window size and hop size of 320 and 80 sample points, respectively. So for the training set, the number of frequency bins $F$ is 513, and the number of temporal frames $T$ is 101.
For the generator, the number of the ConvNeXt block $N$ was set to 8. The period $p$ for each sub-discriminator in the MPD was configured as 2, 3, 5, 7, and 11. In the case of MRAD and MRPD, the FFT point numbers, rectangular window sizes, and hop sizes of the STFT parameter sets were set to [512, 128, 512], [1024, 256, 1024], and [2048, 512, 2048] for the three sub-discriminators, respectively. Both the generator and discriminator were trained until 500k steps using the AdamW optimizer [59], with $\beta_{1}=0.8$, $\beta_{2}=0.99$, and weight decay $\lambda=0.01$. The learning rate was set initially to $2\times 10^{-4}$ and scheduled to decay with a factor of 0.999 at every epoch.  [^2] Audio samples can be accessed at https://yxlu-0102.github.io/AP-BWE, and the source codes will be released after paper acceptance.
@@ IV-C1 Metrics on Speech Quality
  • Virtual Speech Quality Objective Listener (ViSQOL): To access the overall perceived audio quality of the extended speech signals, we employed the ViSQOL  [^3] https://github.com/google/visqol. [60] which uses a spectral-temporal measure of similarity between a reference and a test speech signal to produce a mean opinion score - listening quality objective (MOS-LQO) score. For the audio mode of ViSQOL at a required sampling rate of 48 kHz, the MOS-LQO score ranges from 1 to 4.75, the higher the better. For the speech mode of ViSQOL at a required sampling rate of 16 kHz, the MOS-LQO score ranges from 1 to 5.
@@ IV-C2 Metrics on Generation Efficiency
We used the real-time factor (RTF) as another key metric to evaluate the efficiency of the inference process, which is defined as the ratio of the total inference time for a model processing narrowband input signals to wideband output signals, to the total duration of the wideband signals. In our implementation, the RTFs were calculated using the complete test set on an RTX 4090 GPU and an Intel(R) Xeon(R) Silver 4310 CPU (2.10 GHz).
@@ IV-C3 Metrics on Speech Intelligibility
The main frequency components of human speech are concentrated within the range of approximately 300 Hz to 3400 Hz. This frequency range encompasses crucial information for vowels and consonants, significantly impacting speech intelligibility. Consequently, we analyzed the intelligibility of waveforms extended by speech BWE methods with the target sampling rate of 16 kHz. Firstly, we employed an ASR tool from the ESPNet model zoo  [^4] https://github.com/espnet/espnet_model_zoo. to transcribe the extended 16 kHz speech signals into corresponding texts. Subsequently, we calculated the word error rate (WER) and character error rate (CER) based on the recognition results. Additionally, short-time objective intelligibility (STOI) was also included as an objective metric to indicate the percentage of speech signals that are correctly understood.
TABLE I: Experimental Results in Speech Quality (LSD and ViSQOL) and Generation Efficiency (RTF) for BWE Methods Evaluated on the VCTK Dataset with Target Sampling Rate of 16 kHz, Where in RTF ($a\times$) Representing $a$ Times Real-Time
Method $\mathrm{8kHz\rightarrow 16kHz}$ $\mathrm{4kHz\rightarrow 16kHz}$ $\mathrm{2kHz\rightarrow 16kHz}$ $\mathrm{RTF(CPU)}$ $\mathrm{RTF(GPU)}$
@@
sinc 1.80 4.34 2.68 3.52 3.15 2.73 - -
TFiLM [21] 1.31 4.46 1.65 3.84 1.97 3.10 0.3287 (3.04$\times$) 0.0244 (41.01$\times$)
AFiLM [22] 1.24 4.39 1.63 3.83 1.79 2.75 0.5029 (1.99$\times$) 0.0477 (20.96$\times$)
NVSR [31] 0.79 4.52 0.95 4.11 1.10 3.41 0.7577 (1.32$\times$) 0.0512 (19.54$\times$)
AERO [32] 0.87 4.57 1.00 4.19 - - 0.4395 (2.28$\times$) 0.0217 (46.01$\times$)
AP-BWE* 0.71 4.66 0.88 4.28 0.99 3.77 0.0338 (29.61$\times$) 0.0026 (382.56$\times$)
@@ V-A1 Baseline Methods
For BWE targeting a 16 kHz sampling rate, we first used the sinc filter interpolation as the lower-bound method, and further compared our proposed AP-BWE with two waveform-based methods (TFiLM [21] and AFiLM [22]), a vocoder-based method (NVSR [31]), and a complex-spectrum-based method (AERO [32]). For TFiLM and AFiLM, we used their official implementations  [^5] https://github.com/ncarraz/AFILM.. However, their original papers used the old-version VCTK dataset [61] and employed subsampling to obtain the narrowband waveforms, which aliased high-frequency components. Thus, they performed not a strict BWE task but an SR task. For a fair comparison, we re-trained the TFiLM and AFiLM models with our data-preprocessing manner on the VCTK-0.92 dataset until 50 epochs. For AERO and NVSR, we used their official implementations  [^6] https://github.com/haoheliu/ssr_eval.  [^7] https://github.com/slp-rl/aero.. Notably, AERO did not conduct the experiment at a 2 kHz input sampling rate, and thus, this result was excluded from our analysis.
Additionally, considering some recent BWE methods [31, 47, 48] demonstrated the ability to handle various input sampling rates with a single model, we also trained our AP-BWE with the input sampling rate uniformly sampled from 2 kHz to 8 kHz, denoted as AP-BWE*, to unifiedly extend speech signals at all these three sampling rates to 16 kHz.
@@
TABLE III: Experimental Results in Intelligibility for BWE methods Evaluated on the VCTK Dataset with Target Sampling Rate of 16 kHz

### v2

@@ Abstract
Speech bandwidth extension (BWE) refers to widening the frequency bandwidth range of speech signals, enhancing the speech quality towards brighter and fuller. This paper proposes a generative adversarial network (GAN) based BWE model with parallel prediction of Amplitude and Phase spectra, named AP-BWE, which achieves both high-quality and efficient wideband speech waveform generation. The proposed AP-BWE generator is entirely based on convolutional neural networks (CNNs). It features a dual-stream architecture with mutual interaction, where the amplitude stream and the phase stream communicate with each other and respectively extend the high-frequency components from the source narrowband amplitude and phase spectra. To improve the naturalness of the extended speech signals, we employ a multi-period discriminator at the waveform level and design a pair of multi-resolution amplitude and phase discriminators at the spectral level, respectively. Experimental results demonstrate that our proposed AP-BWE achieves state-of-the-art performance in terms of speech quality for BWE tasks targeting sampling rates of both 16 kHz and 48 kHz. In terms of generation efficiency, due to the all-convolutional architecture and all-frame-level operations, the proposed AP-BWE can generate 48 kHz waveform samples 292.3 times faster than real-time on a single RTX 4090 GPU and 18.1 times faster than real-time on a single CPU. Notably, to our knowledge, AP-BWE is the first to achieve the direct extension of the high-frequency phase spectrum, which is beneficial for improving the effectiveness of existing BWE methods.
@@ I Introduction
In the time domain, speech BWE is a conditionally stringent form of speech super-resolution (SR). Speech SR aims to increase the temporal resolution of low-resolution speech signals by generating high-frequency components, whereas low-resolution speech signals may contain aliased high-frequency components. In contrast, in BWE, only the low-frequency components are preserved in the narrowband signals. Consequently, the BWE task poses greater challenges than SR. Nevertheless, the majority of SR methods are applicable to the BWE task.
@@ III-A1 Generator
where $\bm{\hat{Y}}_{r}=\bm{\hat{Y}}_{a}\cdot\cos(\bm{\hat{Y}}_{p})\in\mathbb{R}^{T\times F}$ and $\bm{\hat{Y}}_{i}=\bm{\hat{Y}}_{a}\cdot\sin(\bm{\hat{Y}}_{p})\in\mathbb{R}^{T\times F}$ denote the real and imaginary parts of the extended short-time complex spectrum, respectively.
@@ III-A2 Discriminator
Directly predicting amplitude and phase and then reconstructing the speech waveform through iSTFT can result in over-smoothed spectral parameters, manifesting as a robotic or muffled quality in the reconstructed waveforms. To this end, we utilize discriminators defined both in the spectral domain and time domain to guide generator $G$ in generating spectra and waveforms that closely resemble real ones. Firstly, considering that the speech signal is composed of sinusoidal signals with various frequencies, with some frequency bands generated through BWE. Due to the statistical characteristics of speech signals varying in different frequency bands, we employ an MPD [40] to capture periodic patterns, with the aim of matching the natural wideband speech across multiple frequency bands. Moreover, since the statistical characteristics of amplitude and phase also differ across frequency bands, and the sole utilization of MPD cannot cover all frequency bands, we consequently define discriminators on both amplitude and phase spectra. Drawing inspiration from the multi-resolution discriminator [41], we respectively introduce MRAD and MRPD, with the aim to capture full-band amplitude and phase patterns at various resolutions. The details of MPD, MRAD, and MRPD are described as follows.
@@ III-B3 Final Loss
$\mathcal{L}_{G}=\sum_{k=1}^{K}\bigg{[}\lambda_{adv}\mathcal{L}_{adv}(G;D_{k})+\lambda_{FM}\mathcal{L}_{FM}(G;D_{k})\bigg{]}+\lambda_{S}\mathcal{L}_{S},$ (13)
@@
where $K$ denotes the numbers of sub-discriminators, and $D_{k}$ denotes the $k$-th sub-discriminator in MPD, MRAD, and MRPD. $\lambda_{adv}$, $\lambda_{FM}$, and $\lambda_{S}$ are hyper-parameters and in all our experiments, we set $\lambda_{S}=1$. For MPD, we set $\lambda_{adv}=1$ and $\lambda_{FM}=1$, while for MRAD and MRPD, we set $\lambda_{adv}=0.1$ and $\lambda_{FM}=0.1$.
@@ IV-B Model Details
We used the same configuration for experiments with target sampling rates of 16 kHz and 48 kHz. For training our proposed AP-BWE model, all the audio clips underwent silence trimming with VCTK silence labels  [^1] https://github.com/nii-yamagishilab/vctk-silence-labels. and sliced into 8000-sample-point segments. To extract the amplitude and phase spectra from raw waveforms, we used STFT with the FFT point number of 1024, Hanning window size, and hop size of 320 and 80 sample points, respectively. So for the training set, the number of frequency bins $F$ is 513, and the number of temporal frames $T$ is 101.
For the generator, the number of the ConvNeXt block $N$ was set to 8. The period $p$ for each sub-discriminator in the MPD was configured as 2, 3, 5, 7, and 11. In the case of MRAD and MRPD, the FFT point numbers, rectangular window sizes, and hop sizes of the STFT parameter sets were set to [512, 128, 512], [1024, 256, 1024], and [2048, 512, 2048] for the three sub-discriminators, respectively. Both the generator and discriminator were trained until 500k steps using the AdamW optimizer [59], with $\beta_{1}=0.8$, $\beta_{2}=0.99$, and weight decay $\lambda=0.01$. The learning rate was set initially to $2\times 10^{-4}$ and scheduled to decay with a factor of 0.999 at every epoch.  [^2] Source codes and audio samples of the proposed AP-BWE can be accessed at https://github.com/yxlu-0102/AP-BWE.
@@ IV-C1 Metrics on Speech Quality
  • Virtual Speech Quality Objective Listener (ViSQOL): To access the overall perceived audio quality of the extended speech signals in an objective manner, we employed the ViSQOL  [^3] https://github.com/google/visqol. [60] which uses a spectral-temporal measure of similarity between a reference and a test speech signal to produce a mean opinion score - listening quality objective (MOS-LQO) score. For the audio mode of ViSQOL at a required sampling rate of 48 kHz, the MOS-LQO score ranges from 1 to 4.75, the higher the better. For the speech mode of ViSQOL at a required sampling rate of 16 kHz, the MOS-LQO score ranges from 1 to 5.
  • Mean Opinion Score (MOS): To further subjectively access the overall audio quality, MOS tests were conducted to evaluate the naturalness of the wideband speech and speech waveforms extended by the speech BWE models. Defining the extension ratio as the ratio between the target sampling rate and the source sampling rate, we selected configurations with the highest extension ratios for subjective evaluations. In each MOS test, twenty utterances from the test set were evaluated by at least 30 native English listeners on the crowd-sourcing platform Amazon Mechanical Turk. For each utterance, listeners were asked to rate a naturalness score between 1 and 5 with an interval of 0.5. All the MOS results were reported with 95% confidence intervals (CI). We also conducted paired $t$-tests to assess the significance of differences between our proposed AP-BWE and the baseline models, reporting $p$-values to indicate the statistical significance of these comparisons.
@@ IV-C2 Metrics on Generation Efficiency
We first used the real-time factor (RTF) to evaluate the inference speed of the model. The RTF is defined as the ratio of the total inference time for processing narrowband source signals into wideband output signals, to the total duration of the wideband signals. In our implementation, RTF was calculated using the complete test set on an RTX 4090 GPU and an Intel(R) Xeon(R) Silver 4310 CPU (2.10 GHz). Additionally, we used floating point operations (FLOPs) to assess the computational complexity of the model. All the FLOPs were calculated using 1-second speech signals as inputs to the models.
@@ IV-C3 Metrics on Speech Intelligibility
The main frequency components of human speech are concentrated within the range of approximately 300 Hz to 3400 Hz. This frequency range encompasses crucial information for vowels and consonants, significantly impacting speech intelligibility. Consequently, we analyzed the intelligibility of waveforms extended by speech BWE methods with the target sampling rate of 16 kHz. Firstly, we employed an advanced ASR model, Whisper [61] to transcribe the extended 16 kHz speech signals into corresponding texts. Subsequently, we calculated the word error rate (WER) and character error rate (CER) based on the transcription results. Additionally, short-time objective intelligibility (STOI) was also included as an objective metric to indicate the percentage of speech signals that are correctly understood.
TABLE I: Experimental Results in Speech Quality (LSD and ViSQOL) and Generation Efficiency (RTF and FLOPs) for BWE Methods Evaluated on the VCTK Dataset with Target Sampling Rate of 16 kHz, Where in RTF ($a\times$) Representing $a$ Times Real-Time
Method $\mathrm{8kHz\rightarrow 16kHz}$ $\mathrm{4kHz\rightarrow 16kHz}$ $\mathrm{2kHz\rightarrow 16kHz}$ $\mathrm{RTF(CPU)}$ $\mathrm{RTF(GPU)}$ $\mathrm{FLOPs}$
@@
sinc 1.80 4.34 2.68 3.52 3.15 2.73 - - -
TFiLM [21] 1.31 4.46 1.65 3.84 1.97 3.10 0.3287 (3.04$\times$) 0.0244 (41.01$\times$) 232.85G
AFiLM [22] 1.24 4.39 1.63 3.83 1.79 2.75 0.5029 (1.99$\times$) 0.0477 (20.96$\times$) 260.76G
NVSR [31] 0.79 4.52 0.95 4.11 1.10 3.41 0.7577 (1.32$\times$) 0.0512 (19.54$\times$) 34.28G
AERO [32] 0.87 4.57 1.00 4.19 - - 0.4395 (2.28$\times$) 0.0217 (46.01$\times$) 141.77G
AP-BWE* 0.71 4.66 0.88 4.28 0.99 3.77 0.0338 (29.61$\times$) 0.0026 (382.56$\times$) 5.97G
@@ V-A1 Baseline Methods
For BWE targeting a 16 kHz sampling rate, we first used the sinc filter interpolation as the lower-bound method, and further compared our proposed AP-BWE with two waveform-based methods (TFiLM [21] and AFiLM [22]), a vocoder-based method (NVSR [31]), and a complex-spectrum-based method (AERO [32]). For TFiLM and AFiLM, we used their official implementations  [^4] https://github.com/ncarraz/AFILM.. However, their original papers used the old-version VCTK dataset [62] and employed subsampling to obtain the narrowband waveforms, which aliased high-frequency components. Thus, they performed not a strict BWE task but an SR task. For a fair comparison, we re-trained the TFiLM and AFiLM models with our data-preprocessing manner on the VCTK-0.92 dataset until 50 epochs. For AERO and NVSR, we used their official implementations  [^5] https://github.com/haoheliu/ssr_eval.  [^6] https://github.com/slp-rl/aero.. Notably, AERO did not conduct the experiment at a 2 kHz source sampling rate, and thus, this result was excluded from our analysis.
Additionally, considering some recent BWE methods [31, 47, 48] demonstrated the ability to handle various source sampling rates with a single model, we also trained our AP-BWE with the source sampling rate uniformly sampled from 2 kHz to 8 kHz, denoted as AP-BWE*, to unifiedly extend speech signals at all these three sampling rates to 16 kHz.
@@
V-A2 Evaluation on Speech Quality
  • Objective Evaluation: As depicted in Table I, our proposed AP-BWE achieved the best performance in speech quality with all kinds of source sampling rates. Compared to sinc filter interpolation, our proposed AP-BWE exhibited significant improvements of 61.7%, 67.5%, and 68.6% in terms of LSD as well as 8.5%, 22.2%, and 37.7% in terms ViSQOL for source sampling rates of 8 kHz, 4 kHz, and 2 kHz, respectively. With the narrowing of source speech bandwidth, the performance advantage of our proposed AP-BWE became more pronounced, indicating the powerful BWE capability of our model. In general, waveform-based methods (TFiLM and AFiLM) performed less effectively than spectrum-based methods (NVSR, AERO, and our proposed AP-BWE), indicating the importance of capturing time-frequency domain characteristics for the BWE task. Within spectrum-based methods, NVSR, relying on high-frequency mel-spectrogram prediction and vocoder-based waveform reconstruction, demonstrated advantages in the LSD metric assessing the extended amplitude. However, the vocoder-based phase recovery was not as effective as the complex spectrum-based approach, so it lagged behind AERO in the ViSQOL metric assessing overall speech quality. Compared to AERO, our AP-BWE, benefiting from explicit amplitude and phase optimizations, successfully avoided the compensation effects between amplitude and phase and consequently achieved better performance in both spectral and waveform-based metrics. Worth noting is that the unified AP-BWE* model exhibited only a slight decrease in performance compared to AP-BWE, and even achieved the highest ViSQOL score at the source sampling rate of 2 kHz. This indicated that our model exhibited strong adaptability to the source sampling rate.
    The key distinction between our approach and others lay in our implementation of explicit high-frequency phase extension. As illustrated in Table II, our proposed AP-BWE consistently outperformed other baselines across various source sampling rates, demonstrating superior performance in terms of instantaneous phase error and phase continuity along both time and frequency axes. For the AP-BWE*, only slight decreases in the AWPD metrics were observed when the source sampling rate was 8 kHz. Under other source sampling rate conditions, the metrics were the same as AP-BWE, indicating the robustness of our unified model on phase. Remarkably, for other baseline methods, some of their AWPD metrics exhibited degradation compared to those of the source sinc-interpolated waveforms. This suggests a limitation in the effective utilization of low-frequency phase information during the speech BWE process by these baseline methods. Moreover, all methods here directly generated waveforms without substituting the original low-frequency components, so their low-frequency phase might be partially compromised, leading to a significant impact on the quality of the extended speech. This observation underscored the critical importance of precise phase prediction and optimization in the context of BWE tasks, further emphasizing the advantage of our approach.
@@ V-A2 Evaluation on Speech Quality
    TABLE III: MOS Tests Results for BWE Methods with Source Sampling Rate of 2 kHz and Target Sampling Rate of 16 kHz
    Methods MOS (CI)
    sinc 3.34 ($\pm$ 0.09)
    TFiLM [21] 3.41 ($\pm$ 0.09)
    AFiLM [22] 3.50 ($\pm$ 0.08)
    NVSR [31] 3.68 ($\pm$ 0.08)
    AP-BWE 3.93 ($\pm$ 0.07)
    Ground Truth 4.01 ($\pm$ 0.06)
  • Subjective Evaluation: To compare the BWE capabilities of our proposed AP-BWE with those of other baseline models, we conducted MOS tests on natural wideband 16 kHz speech waveforms, as well as on speech waveforms extended by AP-BWE and other baseline methods at a source sampling rate of 2 kHz. The subjective experimental results are presented in Table III. For a more intuitive comparison, we visualized the spectrograms of these speech waveforms, as illustrated in Fig.4. According to the MOS results, our proposed AP-BWE outperformed other baseline models very significantly in terms of subjective quality ($p<0.01$). The MOS of TFiLM and AFiLM showed a slight improvement to that of the sinc filter interpolation, demonstrating their insufficient modeling capability for high-frequency components, particularly in the case of high-frequency unvoiced segments as shown in the in Fig. 4. NVSR achieved a decent MOS score compared to TFiLM and AFiLM but still lagged behind our proposed AP-BWE. We can observe that, compared to the spectrograms of natural wideband speech and AP-BWE-extended speech, the NVSR-extended speech spectrogram exhibited relatively low energy in both high-frequency unvoiced segments (e.g., 0.2 $\sim$ 0.3s) and low-frequency harmonics (e.g., 1.1 $\sim$ 1.5s). As a result, the speech signals extended by NVSR would sound duller, negatively impacting its perceived speech quality. In contrast, our proposed AP-BWE effectively extended more robust harmonic structures, demonstrating their strong modeling capabilities and highlighting the effectiveness of explicit predictions of amplitude and phase spectra.
Figure 4: Spectrogram visualization of the original wideband 16 kHz speech waveform and speech waveforms extended by baseline methods and our proposed AP-BWE from the source sampling rate of 2 kHz.
V-A3 Evaluation on Generation Efficiency
We respectively evaluated the generation efficiency of our proposed AP-BWE as well as other baseline methods as outlined in Table I. Considering the inference speed, since NVSR divided the BWE process into the mel-spectrogram extension stage and vocoder synthesis stage, it lagged far behind other end-to-end methods. For TFiLM and AFiLM, since they both operated on the waveform level and utilized RNNs or self-attention to capture long-term dependencies, their inference speeds were consequently constrained. For AERO, although it and our proposed AP-BWE both operated on the spectral level, the utilization of transformer [43] blocks in multiple layers severely slowed down its inference speed. Nevertheless, our AP-BWE model, based on fully convolutional networks and all-frame-level operations, has achieved an astonishingly high-speed waveform generation (29.61 times real-time speed on CPU and 382.56 times on GPU), far surpassing other baseline methods. Considering the models’ computational complexity, the FLOPs of AP-BWE were at least five times smaller than those of the baseline models, further demonstrating the advantage of our proposed model in generation efficiency.
TABLE IV: Experimental Results in Intelligibility for BWE methods Evaluated on the VCTK Dataset with Target Sampling Rate of 16 kHz